OpenStudy (anonymous):

find the derivative of y= log10 x^2-1/ x

OpenStudy (anonymous):

\[y=\log_{10}(\frac{x^2-1}{x})\]?

OpenStudy (anonymous):

probably simplest to break this up as \[\log(x+1)+\log(x-1)-\log(x)\] and differentiate term by term

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

sorry... my computer is acting really goofy... a lot of lag time..

OpenStudy (anonymous):

so you know that \[\frac{d}{dx}\log_{10}(x)=\frac{1}{x\log(10)}\] and apply this to each term

OpenStudy (anonymous):

that last one means natural log of ten. maybe i should write \[\frac{d}{dx}\log_{10}(x)=\frac{1}{x\ln(10)}\] to make it look like it probably does on the book

OpenStudy (anonymous):

so maybe i am being silly. you could write \[\frac{1}{\ln(10)}(\frac{1}{x+1}+\frac{1}{x-1}-\frac{1}{x})\] or i guess you could write \[\frac{1}{\ln(10)}\times\frac{x}{x^2-1}\]